# Online Hypergraph Matching

### v2

@@ Appendix C Integral upper bound for $k$-uniform hypergraphs
  • The offline nodes are partitioned into $k-1$ blocks: $V=C_{1}\cup\dots\cup C_{k-1}$, where $|C_{i}|=2(k-i)$ for each $i\in\{1,\dots,k-1\}$, meaning that $|V|=k\>(k-1)$.
  • The instance first consists of $k-1$ phases with online nodes $w_{1},\dots,w_{k-1}$ arriving, all of which are incident to $2$ hyperedges. For every $i\in\{1,\dots,k-1\}$, both hyperedges incident to $w_{i}$ are disjoint on the offline nodes $V$ and they will both contain $(k-i)$ nodes from $C_{i}$, as well as $1$ node from each $C_{j}$ for $j\in\{1,\dots,i-1\}$.
@@
    $|C_{j}\setminus V(H_{1})|=k-i\qquad\forall j\in\{1,\dots,i\}.$ (17)
    After a certain phase $i-1\leq k-2$, the two hyperedges incident to $w_{i}$ in the next iteration are then both constructed as follows: take $k-i$ nodes from $C_{i}$ and complete it by arbitrarily picking one node from $C_{j}\setminus V(H_{1})$ for every $j\leq i-1$.
  • After phase $k-1$, we have that $|C_{j}\setminus V(H_{1})|=1$ for every $j\leq k-1$, by invariant (17). The instance now makes one more online node $w_{k}$ incident to one hyperedge arrive, whose offline nodes are $C_{j}\setminus V(H_{1})$ for every $j\in\{1,\dots,k-1\}$. Let us also add this hyperedge to $H_{1}$.
Let us first show that (17) holds by induction. In the first phase, both hyperedges partition $C_{1}$ on the offline nodes, since $|C_{1}|=2(k-1)$. One of them is chosen to enter $H_{1}$, meaning that $C_{1}\setminus V(H_{1})=k-1$ after phase $1$. Let us now fix a phase $i\leq k-1$ and suppose that (17) holds for all previous phases. By construction, $C_{i}$ is completely covered by the two hyperedges arriving at phase $i$, since $|C_{i}|=2(k-i)$ and both of these hyperedges contain $k-i$ nodes from $C_{i}$. One of these hyperedges enters $H_{1}$ at the end of phase $i$, meaning that $|C_{i}\setminus V(H_{1})|=k-i$ indeed holds. For any other $C_{j}$ with $j<i$, note that $|C_{j}\setminus V(H_{1})|=k-i+1$ at the beginning of phase $i$, by induction hypothesis. Both hyperedges coming at phase $i$ intersect $C_{j}$ at two different nodes, one of which enters $V(H_{1})$ by the random choice, meaning that $|C_{j}\setminus V(H_{1})|$ drops by $1$ and equals $k-i$, thus showing (17).
@@
Appendix D Rounding algorithm for online hypergraph $b$-matching
In this section, we consider the online hypergraph $b$-matching problem on $k$-uniform hypergraphs, in which every (offline and online) node $v$ can be matched to at most $b$ hyperedges. We show that, for $b=\Omega(\log k)$, any fractional algorithm can be rounded to a randomized integral algorithm while incurring a small loss in the competitive ratio.
@@ Appendix D Rounding algorithm for online hypergraph $b$-matching
Fix some small $0<\epsilon<\frac{1}{2}$ and initialize two empty sets of hyperedges $S,\mathcal{M}\leftarrow\emptyset$. Upon the arrival of an online vertex $w\in W$ with $\delta(w)\subseteq H$ and $x_{h}\in[0,1]$ for every $h\in\delta(w)$, the rounding algorithm is as follows:
@@
Lemma D.1.
Let $\mathcal{A}$ be a fractional algorithm which is $\rho$-competitive. The randomized integral algorithm $R(\mathcal{A},\epsilon)$ then achieves a competitive ratio of at least $(1-\epsilon)(1-k\exp(-\epsilon^{2}b/3))\cdot\rho$.
@@
Consider an arbitrary node $v\in V\cup W$. To bound the probability that $v$ is matched to more than $b$ hyperedges in $S$, we use a Chernoff bound [11, Corollary 1.10]. Fix a node $v$ and a hyperedge $h$, and let $X_{v,h}=\sum_{h^{\prime}\in\delta(v)\setminus\{h\}}\mathbf{1}_{\{h^{\prime}\in S\}}$. Note that $\mu:=\mathbb{E}[X_{v,h}]\leq(1-\epsilon)b$. We now have:
$\Pr\left[X_{v,h}\geq b\right]\leq\exp\left(-\left(\frac{b-\mu}{\mu}\right)^{2}\mu/3\right)\leq\exp(-\epsilon^{2}b^{2}/3\mu)\leq\exp(-\epsilon^{2}b/3),$
@@
$\mathbb{E}[|\mathcal{M}|]$ $=\sum_{h\in H}\Pr[h\in\mathcal{M}]\geq\sum_{h\in H}\Pr[h\in S]\left(1-\Pr[h\in S\setminus\mathcal{M}]\right)\geq\sum_{h\in H}x^{\prime}_{h}\left(1-k\exp(\epsilon^{2}b/3)\right)$
$\geq\left(1-k\exp(\epsilon^{2}b/3)\right)(1-\epsilon)\sum_{h\in H}x_{h}\geq\left(1-k\exp(\epsilon^{2}b/3)\right)(1-\epsilon)\rho\>\mathsf{OPT_{LP}}.$
If $b=C\cdot\log(k)$ for some $C\geq 1$, then by choosing $\epsilon=\sqrt{3\log(C)/C}$ we get that the competitive ratio is at least $(1-\sqrt{3\log(C)/C})(1-1/C)\rho$. By using the $\Omega(1/\log k)$-competitive fractional algorithm from [5], this gives an $\Omega(1/\log k)$-competitive integral algorithm for this setting.

### v3

@@ Appendix D Integral upper bound for $k$-uniform hypergraphs
  • The offline nodes are partitioned into k1k-1 blocks: $V=C_{1}\cup\dots\cup C_{k-1}$, where $|C_{i}|=2(k-i)$ for each $i\in\{1,\dots,k-1\}$, meaning that |V|=k(k1)|V|=k\>(k-1).
  • The instance first consists of k1k-1 phases with online nodes $w_{1},\dots,w_{k-1}$ arriving, all of which are incident to $2$ hyperedges. For every $i\in\{1,\dots,k-1\}$, both hyperedges incident to $w_{i}$ are disjoint on the offline nodes $V$ and they will both contain (ki)(k-i) nodes from $C_{i}$, as well as $1$ node from each $C_{j}$ for $j\in\{1,\dots,i-1\}$.
@@
[…]
    After a certain phase $i-1\leq k-2$, the two hyperedges incident to $w_{i}$ in the next iteration are then both constructed as follows: take kik-i nodes from $C_{i}$ and complete it by arbitrarily picking one node from $C_{j}\setminus V(H_{1})$ for every $j\leq i-1$.
  • After phase k1k-1, we have that $|C_{j}\setminus V(H_{1})|=1$ for every $j\leq k-1$, by invariant (16). The instance now makes one more online node $w_{k}$ incident to one hyperedge arrive, whose offline nodes are $C_{j}\setminus V(H_{1})$ for every $j\in\{1,\dots,k-1\}$. Let us also add this hyperedge to $H_{1}$.
Let us first show that (16) holds by induction. In the first phase, both hyperedges partition $C_{1}$ on the offline nodes, since $|C_{1}|=2(k-1)$. One of them is chosen to enter $H_{1}$, meaning that $C_{1}\setminus V(H_{1})=k-1$ after phase $1$. Let us now fix a phase $i\leq k-1$ and suppose that (16) holds for all previous phases. By construction, $C_{i}$ is completely covered by the two hyperedges arriving at phase $i$, since $|C_{i}|=2(k-i)$ and both of these hyperedges contain kik-i nodes from $C_{i}$. One of these hyperedges enters $H_{1}$ at the end of phase $i$, meaning that $|C_{i}\setminus V(H_{1})|=k-i$ indeed holds. For any other $C_{j}$ with $j<i$, note that $|C_{j}\setminus V(H_{1})|=k-i+1$ at the beginning of phase $i$, by induction hypothesis. Both hyperedges coming at phase $i$ intersect $C_{j}$ at two different nodes, one of which enters $V(H_{1})$ by the random choice, meaning that $|C_{j}\setminus V(H_{1})|$ drops by $1$ and equals kik-i, thus showing (16).
@@
Appendix E Rounding algorithm for online hypergraph $b$-matching
In this section, we consider the online $b$-matching problem on $k$-uniform hypergraphs, in which every (offline and online) node $v$ can be matched to at most $b$ hyperedges. We show that, for $b=\Omega(\log k)$, any fractional algorithm can be converted to a randomized integral algorithm while incurring a small loss in the competitive ratio.
@@ Appendix E Rounding algorithm for online hypergraph $b$-matching
Fix some small $0<\epsilon<1$ and initialize two empty sets of hyperedges $S,\mathcal{M}\leftarrow\emptyset$. Upon the arrival of an online vertex $w\in W$ with $\delta(w)\subseteq H$ and $x_{h}\in[0,1]$ for every $h\in\delta(w)$, the rounding algorithm is as follows:
@@
Lemma E.1.
Let $\mathcal{A}$ be a $\rho$-competitive fractional algorithm. The randomized integral algorithm $R(\mathcal{A},\epsilon)$ achieves a competitive ratio of at least $(1-\epsilon)(1-k\exp(-\epsilon^{2}b/3))\cdot\rho$.
@@
Consider an arbitrary node $v\in V\cup W$. To bound the probability that $v$ is matched to more than $b$ hyperedges in $S$, we use a Chernoff bound [11, Theorem 1.10.1]. Fix a node $v$ and a hyperedge $h$, and let $X_{v,h}=\sum_{h^{\prime}\in\delta(v)\setminus\{h\}}\mathbf{1}_{\{h^{\prime}\in S\}}$. Note that $\mu:=\mathbb{E}[X_{v,h}]\leq(1-\epsilon)b$. If $\mu=0$, it is clear that $x_{v,h}\leq b$, so assume $\mu>0$. We now have:
$\Pr\left[X_{v,h}\geq b\right]$ $\leq\exp\left(-\min\left(\left(\frac{b-\mu}{\mu}\right)^{2},\frac{b-\mu}{\mu}\right)\mu/3\right)$
$\leq\exp\left(-\min\left(\frac{(b-\mu)^{2}}{\mu},b-\mu\right)/3\right)$
$\leq\exp\left(-\min\left(\epsilon^{2}b,\epsilon b\right)/3\right)\leq\exp(-\epsilon^{2}b/3),$
@@
$\mathbb{E}[|\mathcal{M}|]=\sum_{h\in H}\Pr[h\in\mathcal{M}]$ $=\sum_{h\in H}\Pr[h\in S]\left(1-\Pr[h\in S\setminus\mathcal{M}\mid h\in S]\right)$
$\geq\sum_{h\in H}x^{\prime}_{h}\left(1-k\exp(-\epsilon^{2}b/3)\right)$
$=\left(1-k\exp(-\epsilon^{2}b/3)\right)(1-\epsilon)\sum_{h\in H}x_{h}$
$\geq\left(1-k\exp(-\epsilon^{2}b/3)\right)(1-\epsilon)\rho\>\mathsf{OPT_{LP}}.$
If $b=C\cdot\log(k)$ for some $C>6$, then by choosing $\epsilon=\sqrt{6/C}$ we get that the competitive ratio is at least $(1-\sqrt{6/C})(1-\frac{1}{k})\rho$. By using the $\Omega(1/\log k)$-competitive fractional algorithm from [5], this gives an $\Omega(1/\log k)$-competitive integral algorithm for this setting.